OpenStudy (anonymous):

Solve the linear system by the method of elimination. x+4y-z=12 3x+8y-2z=4

jhonyy9 (jhonyy9):

multiply first equation by -3 and assum by second -3x-12y+3z=-36 3x+8y-2z=4 ----------------- 0 -4y +z=-32

jhonyy9 (jhonyy9):

multiply first equation by -2 and assum with second -2x-8y+2z=-24 3x+8y-2z=4 ----------------- x 0 0 =-20 x=-20 -20+4y-z=12 -60+8y-2z=4 4y-z=32 8y-2z=64 / divide by 2 4y-z=32 4y-z=32 -------- 8y-2z=64

OpenStudy (anonymous):

Thanks so much for your help!